### v1

# Towards Long-Term User Welfare in Recommender Systems via Creator-Oriented Information Revelation

Abstract.
Improving the long-term user welfare (e.g., sustained user engagement) has become a central objective of recommender systems (RS). In real-world platforms, the creation behaviors of content creators plays a crucial role in shaping long-term welfare beyond short-term recommendation accuracy, making the effective steering of creator behavior essential to foster a healthier RS ecosystem. Existing works typically rely on re-ranking algorithms that heuristically adjust item exposure to steer creators’ behavior. However, when embedded within recommendation pipelines, such a strategy often conflicts with the short-term objective of improving recommendation accuracy, leading to performance degradation and suboptimal long-term welfare.
The well-established economics studies offer us valuable insights for an alternative approach without relying on recommendation algorithmic design: revealing information from an information-rich party (sender) to a less-informed party (receiver) can effectively change the receiver’s beliefs and steer their behavior. Inspired by this idea, we propose an information-revealing framework, named Long-term Welfare Optimization via Information Revelation (LoRe). In this framework, we utilize a classical information revelation method (i.e., Bayesian persuasion) to map the stakeholders in RS, treating the platform as the sender and creators as the receivers. To address the challenge posed by the unrealistic assumption of traditional economic methods, we formulate the process of information revelation as a Markov Decision Process (MDP) and propose a learning algorithm trained and inferred in environments with boundedly rational creators. Extensive experiments on two real-world RS datasets and across different creator modeling approaches demonstrate that our method can effectively outperform existing fair re-ranking methods and information revealing strategies in improving long-term user welfare.
Recommender System, Information Asymmetry, Information Revelation, Bounded Rationality
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
1. Introduction
With the rapid growth of user-generated content (UGC) on the Web (Krumm et al., 2008), modern recommender platforms (e.g., TikTok, YouTube) have evolved into complex ecosystems where users and content creators continuously interact and influence each other (Bobadilla et al., 2013; Ye et al., 2025). In such dynamic environments, platforms are now shifting their focus from solely optimizing short-term objectives, such as maximizing users’ immediate engagement behaviors (e.g., clicks or purchases), to fostering long-term user welfare (i.e., sustained user engagement), aiming to ensure the lasting health and sustainability of the ecosystem (Xu et al., 2023; Guo et al., 2023).
To enhance the long-term user welfare of recommender systems (RS), previous works have shown that the creators’ creation behavior (i.e., creating new items) is a key factor (Ye et al., 2025; Xu et al., 2023; Yao et al., 2024) beyond recommendation accuracy (i.e., prediction accuracy of user behavior). This is because creators continuously upload new items based on previous user feedback, which shapes the item pool of RS and indirectly influences what users are exposed to in the long run (Ben-Porat and Tennenholtz, 2018; Ben-Porat et al., 2019; Boutilier et al., 2024). Considering the importance of creators’ behavior, most existing works utilize heuristic re-ranking algorithms to steer creators’ content creation. For example, they increase item exposure for small creators to enhance creator retention and stimulate content creation (Guo et al., 2023; Xu et al., 2023; Ye et al., 2024). However, previous works have shown that such a re-ranking strategy, when embedded within recommendation pipelines, often conflicts with the short-term objectives of improving accuracy, inevitably trading off recommendation performance and harming users’ short-term experience (Ye et al., 2024; Xu et al., 2023). Therefore, how to steer creators’ content creation behavior beyond relying on recommendation algorithmic design to enhance the long-term welfare remains a significant challenge in RS.
The investigation of behavioral steering among strategic agents has a long history in economics (Milgrom and Roberts, 1987; Dierkens, 1991; Mishra et al., 1998; Crawford et al., 2018), offering valuable insights for enhancing the long-term welfare of RS. In economics studies, a common approach is to leverage information revelation under conditions of information asymmetry (i.e., agents possess unequal amounts of information) to influence agents’ behavior. Such a strategy aims to design an information revelation strategy, which specifies what information the information-rich party (sender) should disclose to the less-informed party (receiver). By revealing certain information unavailable to the receiver, the receiver can update their beliefs and adjust their strategy accordingly, leading to observable behavioral changes.
Inspired by this idea, we can steer creators’ behavior to enhance long-term welfare by revealing information, viewing the platform as the sender and creators as the receivers. Such a strategy is feasible and practical in real-world platforms, because information asymmetry naturally exists between the platform and creators in RS (Prasad et al., 2023; Boutilier et al., 2024; Ye et al., 2025). Similar but simpler idea have been adopted in industrial practice, offering real-world scenarios and potential applications for more advanced information revelation approaches. For example, platforms like TikTok provide creators with suggestions to produce trending content (see Appendix A).
However, directly applying information revelation to RS is challenging. Traditional methods in economics studies (Yao et al., 2024; Prasad et al., 2023; Boutilier et al., 2024) typically assume that receivers are completely rational, i.e., they can always select the optimal strategy to maximize their utility given the available information. In contrast, real-world creators are bounded rational (Simon, 1990; Kahneman and Tversky, 2013), because their decision-making is constrained by the limited cognitive capacity, attention, and information-processing ability of humans.
To address this challenge, we propose our information-revealing framework tailored for RS, named Long-term Welfare Optimization via Information Revelation (LoRe). Specifically, we map the key components of classical information revelation methods in economics (i.e., Bayesian persuasion (Gan et al., 2022a; Kamenica and Gentzkow, 2011)) onto the corresponding stakeholders in RS. Based on this mapping, we formulate the platform’s information revelation process as a Markov Decision Process (MDP) (Puterman, 1990) and propose a learning algorithm to optimize the information revelation strategy. Instead of relying on the assumption of completely rational agents, LoRe is trained and inferred within a realistic reinforcement learning (RL) environment, where bounded rational users and creators interact with each other.
Unlike previous works that rely on heuristic exposure allocation in the re-ranking stage to steer creators’ behavior (Guo et al., 2023), LoRe serves as an independent module for the platform and does not depend on any specific recommendation algorithm. Experiments on two real-world recommendation datasets demonstrate that LoRe outperforms both creator-fair re-ranking methods and simple information-revealing strategies in improving long-term user welfare (i.e., sustained user clicks). Moreover, since LoRe operates independently, it can be seamlessly integrated with existing ranking or re-ranking algorithms to further enhance long-term welfare.
We summarize our contributions as follows:
(1) We highlight that steering creators’ behavior through information revealing provides an effective approach to enhancing long-term user welfare beyond relying on recommendation algorithmic design.
(2) Inspired by economic studies, we propose a novel information revelation method, LoRe, which models the revelation process as a Bayesian persuasion problem and learns to optimize the revelation strategy in environments where boundedly rational agents interact.
(3) Extensive experiments on real-world datasets demonstrate that our method can effectively improve the long-term user welfare of RS by steering creators’ content creation.
2. Related Work
Multi-Stakeholder RS. Multi-stakeholder RS have received significant attention in recent years in both academic research and industrial practice (Abdollahpouri et al., 2020; Chen et al., 2025). Unlike traditional RS, which primarily focuses on providing the most suitable items based on each user’s request (He et al., 2017; Rendle et al., 2012), multi-stakeholder RS also takes into account the interests of platforms and creators, aligning with the long-term goals of many multi-stakeholder real-world platforms (e.g., TikTok, YouTube). Most existing studies on multi-stakeholder RS can be broadly categorized into two directions: (1) designing better RS algorithms that aim to achieve multi-stakeholder objectives (e.g., provider fairness (Ye et al., 2024; Xu et al., 2025), platform profit (Pei et al., 2019; Azaria et al., 2013)) while maintaining recommendation accuracy; (2) incorporating creators and platforms into the RS environment to enable more comprehensive evaluation of multi-stakeholder RS. For example, some studies focused on developing re-ranking algorithms that allocate exposure resources fairly among different providers to achieve different provider fairness objectives (e.g., Gini coefficient (Do and Usunier, 2022), Max-Min Fairness (Ye et al., 2024; Xu et al., 2023)). Ye et al. (2025) introduces an LLM-based creator simulator, CreAgent, aiming to achieve a more realistic evaluation of multi-stakeholder RS. However, existing works mainly focus on developing and evaluating RS algorithms, and do not consider another crucial side, i.e., steering creators via creation suggestions to indirectly optimize the long-term welfare of RS.
Information asymmetry in RS. A relevant topic to our work is the literature considering the information asymmetry in RS (Prasad et al., 2023; Yao et al., 2023, 2024; Boutilier et al., 2024; Ye et al., 2025). Boutilier et al. (2024); Ye et al. (2025) highlight the significant information asymmetry in RS, but they do not proposed a method to improve the welfare of RS under such asymmetry. Prasad et al. (2023) propose a prompting policy that promises to provide creators a certain level of user clicks to influence the their content creation. Yao et al. (2023, 2024) suggest influencing content creation by adjusting reward weights to encourage different types of content. However, these studies model creators as completely rational agents, ignoring the bounded rationality observed in real-world creators, which limits the applicability and performance of their methods.
Information revelation under information asymmetry. Our work is also related to the literature on information revelation under information asymmetry in economics, where a prominent approach is Bayesian persuasion (Kamenica and Gentzkow, 2011). In Bayesian persuasion, an informed party (sender) designs an information revelation strategy to reveal the optimal information to a less informed party (receiver) (Kamenica and Gentzkow, 2011). The most relevant Bayesian persuasion models to our work consider one sender (the platform) and multiple receivers (the creators) (Wang, 2013; Arieli and Babichenko, 2019; Castiglioni et al., 2021; Bernasconi et al., 2023; Zhao et al., 2025), where the sender simultaneously reveals information to multiple receivers. These studies investigate how to design information disclosure strategies to maximize the sender’s utility in a static setting. There is also a stream of research on dynamic Bayesian persuasion (Gan et al., 2022b, a; Wu et al., 2022), which examines repeated interactions between a sender and a receiver. However, because of their strong assumption of completely rational receivers, these models cannot be directly applied to real-world RS scenarios, where creators typically exhibit bounded rationality.
3. Preliminaries
In this section, we first introduce the settings of real-world RS. Then we introduce the basic idea of Bayesian Persuasion. Finally, we describe how to model creators as bounded rational agents.
3.1. Recommender Systems
Stakeholders in RS. Considering a real-world content recommendation platform (e.g., YouTube, TikTok). The platform needs to recommend items (e.g., videos, articles, etc.) produced by content creators to users. Let $\mathcal{U}=\{u_{i}\}_{i=1}^{|\mathcal{U}|}$, $\mathcal{C}=\{c_{i}\}_{i=1}^{|\mathcal{C}|}$, $\mathcal{G}=\{g_{i}\}_{i=1}^{|\mathcal{G}|}$ represent the set of users, creators, and item genres that creators can produce. Besides, let $\mathcal{I}$ denote the current corpus of items that can be recommended to users.
Interactions between different stakeholders. We model the interaction between the platform, creators, and users as a dynamic process with total $T$ interaction rounds. At the beginning of each round $t\in\{1,2,\dots,T\}$, each creator $c_{j}$ will produce an item with the genre in $\mathcal{G}$. The items produced by creators in each round will be added to $\mathcal{I}$. Then, some users will arrive and access the platform. For each accessing user, the platform needs to select $k$ items from $\mathcal{I}$ for recommendation. Each user can choose to click or skip the items recommended to him. After that, the platform records the interaction between users and items. Specifically, we use $\bm{Y}$ to represent the click matrix, where $\bm{Y}_{i,j,t}=1$ if item $i\in\mathcal{I}$ is clicked by user $u_{j}$ in round $t$ and 0 otherwise. In addition, we use $\mathcal{H}_{c_{j}}=\{(i,k_{i})|i\in\mathcal{I}\}$ to denote historical information of items created by creator $c_{j}$, where $(i,k_{i})\in\mathcal{H}_{c_{j}}$ means that the item $i\in\mathcal{I}$ produced by creator $c_{j}$ has been clicked $k_{i}$ times. For easy notation, let $\mathcal{H}=\{\mathcal{H}_{c_{1}},\mathcal{H}_{c_{2}},\dots,\mathcal{H}_{c_{|\mathcal{C}|}}\}$ denote the set of all creators’ historical information.
Information asymmetry in RS. In real-world RS, the platform has the historical information $\mathcal{H}$ of all creators and the interaction information $\bm{Y}$ of all users and items. In contrast, each creator $c_{j}$ only has its own historical information $\mathcal{H}_{c_{j}}$. Obviously, there is a significant information asymmetry between platform and creators. Therefore, how to improve the long-term welfare under information asymmetry is an unexplored and urgent problem to be addressed.
Figure 1. The overall framework of our method LoRe, which consists two phases: Online Inference and Offline Training.
3.2. Bayesian Persuasion
Information asymmetry formulation. The standard Bayesian persuasion considers an idealized model with a sender and a receiver. Both the receiver’s utility function $\mu_{r}(a_{r},w)$ and the sender’s utility function $\mu_{s}(a_{r},w)$ depend on receiver’s action $a_{r}\in\mathcal{A}_{r}$ and state $w\in\mathcal{W}$. The state $w$ follows a cumulative distribution function $F(w)$ with probability density function $f(w)$. It is assumed that both sender and receiver know $F(w)$, but the sender exactly knows the real state $w$ after $w$ is sampled. Thus, the sender have more information than the receiver, which causes the information asymmetry.
Bayesian persuasion under information asymmetry. The receiver will choose an optimal action $a_{r}^{*}$ to maximize her expected utility, and the sender will also obtain a corresponding utility. However, the sender can use a so-called signaling scheme to reveal the receiver some information, and then influence receiver’s action. Specifically, we use $\sigma=(\mathcal{A}_{s},\pi)$ to denote a signaling scheme, where $\mathcal{A}_{s}$ is the set of possible signals (messages) and $\pi:\mathcal{W}\times\mathcal{A}_{s}\mapsto[0,1]$ is the signaling strategy which maps the state to a distribution of signals. For instance, $\pi(a_{s}|w)$ denotes the probability of sending signal $a_{s}\in\mathcal{A}_{s}$ when the real state is $w$. According to the revelation principle (Myerson, 1979), each signal $a_{s}$ can be viewed as an action suggestion for receiver, that is $|\mathcal{A}_{r}|=|\mathcal{A}_{s}|$, where $a_{s}\in\mathcal{A}_{s}$ means suggesting receiver taking corresponding action $a_{r}\in\mathcal{A}_{r}$ and receiver will follows sender’s suggestion [^1] Some constraints can be used to ensure that the receiver will follow the sender’s suggestion, and these constraints do not affect the optimal solution. For details, please refer to the reference (Kamenica and Gentzkow, 2011).. By designing the optimal strategy $\pi^{*}$, the sender can improve both the utility of sender and receiver under information asymmetry (Kamenica and Gentzkow, 2011).
Table 1. Correspondence between elements in the Bayesian Persuasion and RS. See Section 4 for undefined notations.
Bayesian Persuasion Recommender System
sender platform
receivers creators
state $w$ content distribution $\bm{e}$ and trust $\bm{d}$
signal set $\mathcal{A}_{s}$ suggestion set $\mathcal{A}_{p}$
signaling strategy $\pi$ information revealing strategy $\pi$
3.3. Bounded Rationality Modeling
Motivated by the advantage of Bayesian persuasion under information asymmetry, we adopt the idea of this framework to guide how the platform reveals information in RS. Simply put, we allow the platform to send suggestions to creators, i.e., suggesting creators creating specific genres of content, which is called the platform’s information revelation strategy.
In the standard Bayesian persuasion, receivers are assumed to be completely rational, which means they can precisely choose an action $a_{r}^{*}$ to maximize their utility $\mu_{r}(a_{r},w)$. This also allows senders to design signaling strategy that ensure that receivers completely follow the sender’s action suggestion. However, creators in reality are typically bounded rational, as they cannot compute a specific utility function to maximize their utility. Thus, platform in RS cannot make sure that creators will follow platform’s suggestion. To model creators’ bounded rationality, inspired by Prasad et al. (2023), we design a trust-based model, where each creator $c_{i}$ has a trust degree $d_{{}_{i}}\in[0,1]$. When creator $c_{i}$ receives platform’s suggestion, the probability that creator $c_{i}$ follows the suggestion is $d_{{}_{i}}$ .
If creator $c_{i}$ dose not follow platform’s suggestion, we assume that he can use any strategy that does not perform utility calculation to choose an action, such as randomly selecting from the genres of previously created content, or selecting the most popular genre of previously created content, etc. That is, we do not restrict the creators’ decision model, and any decision model can be embedded in our approach that is introduced in next section. In the experiment section, to verify the universality of our method for different creator decision models, we use different model to simulate creators’ decision-making, such as LLM-based simulation model and heuristic-based simulation models.
We correspond the elements of RS with those in the Bayesian persuasion, as shown in Table 1.
$\bullet$ The platform and creators are regarded as the sender and receivers, respectively. This is because the platform, who possesses more information, can guide the creator’s creation by revealing information.
$\bullet$ The content distribution in RS and creators’ trust degree are viewed as the state in Bayesian persuasion.
$\bullet$ The suggestion that platform sends to the creators is considered as the signal in Bayesian persuasion.
$\bullet$ The platform’s strategy of sending suggestions (information revelation strategy $\pi$) is corresponds to the signaling strategy $\pi$ in Bayesian persuasion.
4. Our Approach: LoRe
In this section, to take into account creators’ bounded rationality and optimize platform’s information revelation strategy, we propose a novel method, LoRe. The overall algorithm workflow of LoRe is shown in Figure 1 and Algorithm 1. To optimize the long-term welfare of RS, we first formulate the dynamic interaction process under information asymmetry between the platform and creators as a MDP, which is described in Section 4.1. Then, we design a RL algorithm to optimize the platform’s information revelation strategy under such a dynamic environment, as presented in Section 4.2.
4.1. MDP for Platform and Creators Interactions
In our method, we use $\mathcal{A}_{p}=\{a_{i}\}_{i=0}^{|\mathcal{G}|}$ to denote the set of platform’s suggestions, where $a_{0}$ means no suggestion and $a_{i}\ (i\neq 0)$ means that platform suggests the creator to produce item with genre $g_{i}$. At the beginning of each round $t$, the platform will send suggestions selected from $\mathcal{A}_{p}$ to creators. Note that different creators may receive different suggestions. After receiving the suggestion, each creator will decide whether to follow the suggestion and choose to produce an item. Then, for each arriving user, the platform will generate $k$ items to recommend and each user can choose to click or skip these items. Then, the MDP for the interaction between platform and creators can be defined as a tuple $M=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where:
$\bullet$ State $\mathcal{S}$: $\mathcal{S}$ is the set of states. Each state is denoted by
$\bm{s}^{t}=(e_{1}^{t},e_{2}^{t},\dots,e_{|\mathcal{C}|}^{t};\hat{d}_{1}^{t},\hat{d}_{2}^{t},\dots,\hat{d}_{|\mathcal{C}|}^{t})\in\mathcal{S},$
which contains the distribution of item genres and the predicted creators’ trust in round t1t-1, where $e_{i}^{t}$ denotes the genre of item created by creator $c_{i}$ and $\hat{d}_{i}^{t}$ represents the predicted trust of creator $i$ in round t1t-1. We use $\bm{e}=(e_{1},e_{2},\dots,e_{|\mathcal{C}|})$ to denote content distribution and use $\bm{\hat{d}}=(\hat{d}_{1},\hat{d}_{2},\dots,\hat{d}_{|\mathcal{C}|})$ to represent the predicted trust vector. In simple terms, we have $\bm{s}^{t}=(\bm{e}^{t};\bm{d}^{t})$.
$\bullet$ Action $\mathcal{A}$: $\mathcal{A}$ is the set of the platform’s actions. Platform’s each action $\bm{a}^{t}\in\mathcal{A}$ is a vector $(a_{1}^{t},a_{2}^{t},\dots,a_{|\mathcal{C}|}^{t})$, where $a_{i}^{t}\in\mathcal{A}_{p}$ represents the suggestion sent to creator $c_{i}$ in round $t$.
$\bullet$ Transition function $\mathcal{T}$: The transition probability function is defined as $\mathcal{T}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$, where $\mathcal{T}(\bm{s}^{t+1}|\bm{s}^{t},\bm{a}^{t})$ denotes the probability that the state at round $t+1$ is $\bm{s}^{t+1}$ given that the the state at round $t$ is $\bm{s}^{t}$ and the platform takes action $\bm{a}^{t}$.
$\bullet$ Reward function $\mathcal{R}$: The reward function is defined as $\mathcal{R}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathbb{R}$, where $\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})$ denote the total user clicks (the measure of welfare) in round $t$, given that state at round $t$ is $\bm{s}^{t}$, the platform takes action $\bm{a}^{t}$, and the state transitions to $\bm{s}^{t+1}$ in round $t+1$.
$\bullet$ Discount rate $\gamma$: $\gamma\in[0,1]$ is the discount factor that quantifies the present value of future rewards.
Note that we incorporate predicted creators’ trust degree into the state, which can help capture creators’ bounded rational behavior. The platform’s strategy of sending suggestion signals is defined as $\pi:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$, where $\pi(\bm{a}^{t}|\bm{s}^{t})$ is the probability of taking action $\bm{a}^{t}$ when the state at round $t$ is $\bm{s}^{t}$. Then, the expected discounted user clicks under strategy $\pi$ can be formulated as:
(1) $J(\pi)=\operatorname{\mathbb{E}}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathcal{R}(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1})\right],$
where $\tau$ is a trajectory, i.e., $\tau=(\bm{s}^{0},\bm{a}^{0},\bm{s}^{1},\bm{a}^{1},\dots)$ and $\tau\sim\pi$ is a shorthand indicating that the distribution over trajectories depends on $\pi:\bm{a}^{t}\sim\pi(\cdot|\bm{s}^{t})$, $\bm{s}^{t+1}\sim\mathcal{T}(\cdot|\bm{s}^{t},\bm{a}^{t})$. Furthermore, we respectively use $V^{\pi}(\bm{s}^{t})$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ to denote the state value function, state-action value function and the advantage function, where $V^{\pi}(\bm{s}^{t})$ is the value of state $\bm{s}^{t}$, $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is the value of taking action $\bm{a}^{t}$ under state $\bm{s}^{t}$ and $A^{\pi}(\bm{s}^{t},\bm{a}^{t})=Q^{\pi}(\bm{s}^{t},\bm{a}^{t})-V^{\pi}(\bm{a}^{t})$ [^2] In practice, the advantage function $A^{\pi}(\bm{s}^{t},\bm{a}^{t})$ is calculated by the generalized advantage estimation (Schulman et al., 2015).. $V^{\pi}(s^{t})$ and $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$ are common functions in MDP, and their definitions can be found in appendix B.
4.2. Reinforcement Learning Algorithm for Platform’s Strategy
In this section, we aim to optimize the long-term welfare in Eq(1) under the dynamic user-creator-platform interaction environment by learning an information-revealing strategy $\pi$. To achieve this, we design an actor-critic based RL algorithm. The overall workflow of our algorithm is shown in Algorithm 1, which contains two phases: the online inference (Line 4-13) and the offline training (Line 14-19).
0: Discount rate $\gamma$, clip hyperparameter $\epsilon$, learning rate lrlr, total rounds $T$, training epochs $M$, cumulative data rounds NN, reply buffer $\mathcal{B}=\{(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})\}$.
0: The policy loss $L(\pi_{\theta})$ decreases by less than 1% over 10 consecutive rounds.
1: Initialize state $\bm{s}^{1}$ from log data.
2:for $t=1$ to $T/N$ do
3:  Clear reply buffer: $\mathcal{B}=\{\}$.
4:  // $\texttt{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0} Phase 1: Online Inference}}$.
5:  for $r=1$ to NN do
6:   Get item genres’ distribution $\bm{e}^{t}$ and predict trust $\hat{\bm{d}}^{t}$.
7:   Get state $\bm{s}^{t}=(\bm{e}^{t},\hat{\bm{d}}^{t})$.
8:   Sample action $\bm{a}^{t}\sim\pi_{\theta}(\bm{s}^{t})$. $\texttt{Platform sends suggestions}$.
9:   $\texttt{Creators produce items}$.
10:   The next state transitions to $\bm{s}^{t+1}$.
11:   $\texttt{Users interact with items (click or skip)}$.
12:   Get the total users’ clicks $\mathcal{R}$.
13:   Add $(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ to reply buffer $\mathcal{B}$.
14:  end for
15:  // $\texttt{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1} Phase 2: Offline Training}}$.
16:  for $i=1$ to $M$ do
17:   Calculate actor loss $L$ and critic loss $L^{TD}$ based on $\mathcal{B}$ according to Eq. (2) and Eq. (3).
18:   Update actor network using the gradient $\nabla_{\pi_{\theta}}L$.
19:   Update critic network using the gradient $\nabla_{V_{\phi}}L^{TD}$.
20:  end for
21:end for
Algorithm 1 The overall workflow of our method LoRe.
4.2.1. Online Inference
At the beginning, the platform gets the current item genres’ distribution and predicts creators’ trust degree, which together constitute the state $\bm{s}^{t}$. We show how to predict creators’ trust degree in the following subsection. Then, the actor network $\pi_{\theta}$, which is parameterized by $\theta$, takes $\bm{s}^{t}$ as input and outputs the action distribution $\bm{P}^{t}=\pi_{\theta}(\bm{s}^{t})$, where the input $\bm{s}^{t}$ and output $\bm{P}^{t}$ are illustrated as follows.
Input. As shown in Figure 1, we represent the state $\bm{s}^{t}\in\mathbb{R}^{|\mathcal{C}|\times(|\mathcal{G}|+2)}$ as a one-hot matrix, where each row of the matrix is the one-hot vector of the item created by corresponding creator connected with the corresponding trust degree. The length of the one-hot vector is $|\mathcal{G}|+2$, because there are cases where the creator dose not create item.
Output. In the classic actor network, the actor usually outputs a distribution over the action space $\mathcal{A}$. However, the number of possible actions is $|\mathcal{A}|=(|\mathcal{G}|+1)^{|\mathcal{C}|}$, which is too large to handle. To address this problem, we design the actor to output the distribution over suggestion set $\mathcal{A}_{p}$ rather than the distribution over action set $\mathcal{A}$. That is, instead of outputting a joint distribution over all platform’s possible actions, we output an independent distribution over the suggested actions for each creator. Specifically, the output of actor is $\bm{P}=\pi_{\theta}(\bm{s}^{t})\in\mathbb{R}^{|\mathcal{C}|\times(\mathcal{G}|+1)}$, where each row of $\bm{P}$ is a distribution over the suggestion set for corresponding creator. Then, the suggestions sent to creators are sampled from the corresponding row of $\bm{P}$, which constitute the platform’s action $\bm{a}^{t}\in\mathbb{R}^{|\mathcal{C}|}$ as shown in Figure 1.
After sending suggestions to creators, creators produce items and the genres of these new items with creators’ trust degree constitute the next state $\bm{s}^{t+1}$. After that, platform recommends items to each user and get the total users’ clicks $\mathcal{R}$ as reward. Finally, the data $b^{t}=(\bm{s}^{t},\bm{a}^{t},\bm{s}^{t+1},\mathcal{R})$ is then added to the reply buffer $\mathcal{B}$ which is used to train the actor and critic networks in the offline training phase.
4.2.2. Offline Training
We use the collected reply buffer $\mathcal{B}$ to train the actor and critic networks. We denote the critic network by $V_{\phi}(\bm{s})$, which is parameterized by $\phi$. The critic network takes state $\bm{s}^{t}$ as input and outputs the estimated state value of $\bm{s}^{t}$, which is used to approximate the state value function $V^{\pi}(\bm{s})$ and help optimize the actor network.
Actor update. To update the actor, we use the clipped surrogate objective (Schulman et al., 2017) as the loss function:
(2) $L(\pi_{\theta})=-\sum_{b\in\mathcal{B}}\min\left(\pi^{\prime}(\bm{a}|\bm{s}),\text{clip}\left(\pi^{\prime}(\bm{a}|\bm{s}),1-\epsilon,1+\epsilon\right)\right)A^{\pi_{\theta}}(\bm{a},\bm{s}),$
where $\pi^{\prime}(\bm{a}|\bm{s})=\frac{\pi_{\theta}(\bm{a}|\bm{s})}{\pi_{\theta_{old}}(\bm{a}|\bm{s})}$, $\text{clip}(x,l,r)=\max(\min(x,r),l)$ and $\epsilon$ is the clip hyperparameter. $\pi_{\theta_{old}}$ is the actor network after the last update and remains unchanged when the current actor network $\pi_{\theta}$ is updated.
Critic update. To update the critic network, we use the MSE of the temporal-difference error (TD-error) as the loss function:
(3) $L^{TD}(V_{\phi})=\sum_{b\in\mathcal{B}}(\mathcal{R}_{t}+\gamma V_{\phi}(\bm{s}^{t+1})-V_{\phi}(\bm{s}^{t}))^{2}.$
We determine whether our algorithm has converged based on the policy loss $L(\pi_{\theta})$. Specifically, we regard the algorithm as converged when the policy loss remains nearly flat, defined as a decrease of less than 1% over 10 consecutive episodes.
4.2.3. Trust prediction
We use maximum likelihood estimation (MLE) to predict creators’ trust degrees. The training dataset $\mathcal{D}=\{(i,y)\}$ is collected from the interactions between the platform and creators, where $i$ denotes the creator’s ID, and $y\in\{0,1\}$ is a binary label indicating whether creator $i$ follows the platform’s suggestion. A neural network parameterized by $\psi$, denoted as $P_{\psi}$, is employed to model the creators’ trust. The network takes a creator’s ID as input and outputs the estimated trust degree, i.e., $\hat{d}_{i}=P_{\psi}(i)$. The parameters $\psi$ are learned by minimizing the MLE loss:
$L(P_{\psi})=-\sum_{(i,y)\in\mathcal{D}}\left[y\log(P_{\psi}(i))+(1-y)\log(1-P_{\psi}(i))\right].$
Figure 2. Comparison of user clicks in RS between our method and other baselines on two different datasets.
5. Experiments
In this section, we conduct extensive experiments to verify the effectiveness of our method in improving the long-term welfare of RS by addressing the following research questions:
RQ1: Can our method improve the long-term welfare of RS?
RQ2: What is the inherent insight that our method can enchance the long-term welfare?
RQ3: How robust our method is, and whether it can perform well under various settings?
The source code of LoRe and datasets are available  [^3] https://anonymous.4open.science/r/LoRe-E792.
5.1. Experimental Setups
Real-world dataset. We conduct experiments on two large-scale recommendation datasets, both collected from real-world recommendation platforms. YouTube is a dataset crawled from the world’s leading recommendation platform (Ye et al., 2025). It contains rich information about content creators across diverse categories, along with user comments on their uploaded items. The dataset comprises 4,004 creators, 1.97 million users, and 0.19 million distinct items spanning 14 genres. Amazon Video Games (Amazon)  [^4] https://amazon-reviews-2023.github.io/ is a widely used recommendation dataset. It consists of user-generated reviews, ratings, and product metadata collected from the Amazon e-commerce platform. It contains 2.8M users, 137.2K items, and 1326 creators.
Simulation setups. To better approximate real-world RS environments, we adopt the multi-stakeholder RS simulation platform proposed by Ye et al. (2025), which integrates LLM-based agents to simulate the dynamic behaviors of both users and creators. Recent studies have demonstrated that LLMs are capable of accurately mimicking user preferences and the decision-making of bounded-rational creators (Wang et al., 2024; Zhang et al., 2024; Wang et al., 2025; Ye et al., 2025). To test the robustness and generalization ability of our method under diverse environments, we also conduct additional experiments under environments where creators are simulated by non-LLM-based models (See Section 5.4). Besides, we consider the active behaviors of creators and users, as well as creator retention, following previous works (Mladenov et al., 2020; Xu et al., 2023; Ye et al., 2025): each creator $c_{i}\in\mathcal{C}$ and each user $u_{i}\in\mathcal{U}$ has probability $p_{c_{i}}$ and $p_{u_{i}}$ to be active, respectively. In each round, these probabilities determine whether a creator or user is active. Only active creators create items, and only active users receive recommendations. Moreover, a creator $c_{i}$ is assumed to leave the platform if no clicks are received in $l$ consecutive rounds of creation.
Table 2. The content diversity (CD) and average number (AN) of active creators across different methods on different datasets.
YouTube Amazon
Method CD AN CD AN
LoRe (ours) 3.259 16.250 3.462 14.710
Reweight 3.223 15.625 2.744 10.530
P-MMF 3.196 15.930 2.697 11.025
FairRec 3.182 14.245 2.646 9.880
Most History Click 3.197 14.730 2.805 11.610
Most Click 3.186 13.840 2.847 11.775
w/o signal 3.177 11.750 2.546 9.665
Baselines. We compared our method with several baselines, including: w/o Signal: revealing no information to the creator at all, except for their own creation history; and some naive information revelation strategies: Most History Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres the creator has ever created; Most Click: suggesting each creator to produce the item genre with the highest number of user clicks across all genres. In addition, we compared our method with approaches that focus on recommendation algorithm design, such as P-MMF (Xu et al., 2023) and FairRec (Guo et al., 2023). We also included a baseline that influences creators by adjusting their rewards, namely Reweight (Yao et al., 2024).
Metrics. Following previous works (Ye et al., 2025; Yao et al., 2024), we use the cumulative user click number to measure the long-term welfare. We adopt RecAgent (Wang et al., 2025) as the user simulator, where users can choose to click or skip when presented with an item. The number of clicks by all users in each round is recorded as the user welfare for that round.
Figure 3. The number of genres created by each creator.
5.2. Effectiveness of LoRe (RQ1)
To evaluate the effectiveness of LoRe in enhancing long-term user welfare, we compare its performance against several baselines on two recommendation datasets. In this experimental setting, the number of creators is fixed at 50 and the number of users at 100. The platform employs DIN (Zhou et al., 2018) as the based recommendation model, and both user and creator agents are powered by LLaMA3-8B (Touvron et al., 2023). In addition, we initialize creators’ trust degrees using a truncated Gaussian distribution, reflecting real-world scenarios where agents exhibit moderate trust levels, and extremely high or low trust is uncommon (Teacy, 2006; Teacy et al., 2012). Denoted by $\mathcal{N}(\mu,\sigma;0,1)$ a truncated Gaussian distribution with mean $\mu$, variance $\sigma$, lower bound 0, and upper bound $1$. Then, each creator’s trust degree is randomly sampled from distribution $\mathcal{N}(0.5,1.0;0,1)$.
Figure 2 shows the changes in cumulative user clicks per round as the number of rounds increases. LoRe consistently achieves higher user welfare than all baselines on both the YouTube and Amazon datasets. Compared with approaches that only adjust recommendation algorithms or creator rewards (e.g., PMMF, FairRec, Reweight), LoRe performs substantially better, indicating that influencing creator behavior through selective information revelation is more effective for improving long-term user welfare.
Existing methods also show inconsistent performance across datasets. For instance, Reweight outperforms all other baselines on YouTube but performs worse than naive information revelation strategies on Amazon. This is primarily due to the more skewed user distribution in the Amazon dataset compared to the YouTube dataset, i.e., the majority of users in the Amazon dataset favor a particular genre of content, while fewer users favor other genres, which means that existing methods may significantly degrade in performance under such conditions.
Naive information revelation strategies (Most History Click, Most Click) also display unstable performance: they outperform w/o Signal on Amazon but are similar to it on YouTube. This suggests that naive strategies lack robustness across recommendation environments. In contrast, LoRe consistently achieves the best long-term welfare, demonstrating that strategic information revelation is not only more effective than naive approaches but also more robust across different environments.
Figure 4. Training time cost and the influence of trust degree.
5.3. Inherent insight of LoRe (RQ2)
To better understand the inherent insight of LoRe, we analyze the dynamics of RS from both macro and micro perspectives. At the macro level, we examine the content diversity and the average number of active creators during the interactions among the platform, creators, and users. At the micro level, we investigate the variation in the number of content genres produced by individual creators under different methods.
Figure 5. The performance of different methods under various creator simulation models.
Macro perspective. To evaluation the content diversity, we employ entropy which is a widely used metric for measuring diversity (Spellerberg and Fedor, 2003). Let $p$ denote the probability density function over genre set $\mathcal{G}$, where $p(g_{i})\ (g_{i}\in\mathcal{G})$ represents the probability of content genre $g_{i}$. Then, the content diversity is defined as $D(p)=\sum_{g_{i}\in\mathcal{G}}-p(g_{i})\log p(g_{i})$. Diversity is an important property of RS, and increasing diversity is often associated with improved user welfare (Castells et al., 2021). Intuitively, higher diversity reflects a more balanced distribution of content genres, whereas lower diversity corresponds to a skewed content distribution in which most content is concentrated in only a few categories. Such extreme content distribution may prevent some users from accessing their preferred content, resulting in user churn and reduced user welfare. Therefore, greater diversity is more conducive to improving users’ long-term welfare. As shown in Table 2, our method achieves the highest content diversity across different datasets. This demonstrates that our approach enhances the diversity of RS compared with other baselines, which in turn indirectly improves user welfare.
In our simulations, we assumed that a creator would exit after receiving no clicks for 10 consecutive rounds. We calculated the average number of active creators under different methods across datasets. As shown in Table 2, our method retains more creators than other baselines. This improvement arises because the platform can influence creators’ behavior by suggesting them creating specific content genre, thereby guiding them to produce content that attracts more clicks.
Micro perspective. We further analyze the variation in the number of content genres produced by each creator. As shown in Figure 3, the x-axis represents the creator ID, and the y-axis indicates the number of distinct genres created by each individual. Obviously, compared to the baseline "w/o Signal", our approach leads to an increase in the number of genres per creator, which directly contributes to the overall improvement in content diversity.
5.4. Robustness Analysis (RQ3)
In this section, we conduct extensive experiments to evaluate the robustness of our method.
Training cost. We evaluate the convergence rounds and net training time of our algorithm at different scales to demonstrate its scalability in large-scale simulations. As shown in Figure 4 (a), our algorithm requires only a few rounds to converge. When the number of agents is 50, it converges in fewer than 70 rounds; when the number of agents increases to 1,000, convergence is achieved in fewer than 40 rounds. In addition, the net training time per round remains within an acceptable range. Even with 1,000 agents, the net training time per round is less than 4 seconds. These results confirm that our method scales effectively to large-scale, real-world scenarios.
Influence of creator simulation models. As discussed in Section 3.3, our method is capable of handling various creator simulation models. To verify this and evaluate its robustness, we conduct experiments across different creator simulation models. Specifically, we consider both LLM-based simulation models: CreAgent (Ye et al., 2025) (with Mistral-7B and Qwen2.5-7B as backbones), and heuristic-based models: Creator Feature Dynamics (CFD) (Lin et al., 2024; Zhan et al., 2021), where creators adjust their creation strategies using user feedback as gradients scaled by a learning rate, and SimuLine (Zhang et al., 2023), where creators’ subsequent creations are determined by probabilistic sampling based on the number of clicks in previous steps.
Figure 5 presents the performance comparison between our method and the baselines under different creator simulation models. The results clearly show that, regardless of whether the simulation model is LLM-based or heuristic-based, our approach consistently outperforms all baselines. In contrast, existing recommendation-algorithm-based methods (i.e., P-MMF and FairRec) exhibit inconsistent performance across different creator simulation models. For instance, FairRec performs best among baselines under the Mistral-7B, Qwen2.5-7B, and SimuLine simulation models but is outperformed by P-MMF under the CFD model. Moreover, Reweight, which influences creator behavior by adjusting their rewards, performs relatively poorly. This is because it is designed for a specific creator decision model and thus may fail to generalize well when the decision model changes.
Table 3. The average user clicks achieved through combining LoRe with FairRec or P-MMF on YouTube and Amazon datasets.
Method YouTube Amazon
LoRe (ours) 137.4954 95.1499
P-MMF 108.5688 51.2752
FairRec 113.2789 46.1865
LoRe + P-MMF 141.2263 101.9969
LoRe + FairRec 146.0092 98.6177
Influence of creators’ trust degree. Although platforms can provide various suggestions to creators, real-world creators differ in their degree of trust toward the platform. Highly trusting creators are more likely to follow the platform’s suggestions, whereas those with low trust are less likely to be influenced. In addition, in practice, creators’ trust levels may evolve over time according to the number of clicks they receive Prasad et al. (2023).
To evaluate the effectiveness of our method under different levels of creator trust, we conduct experiments where creators’ trust degrees are sampled from distributions with varying expectations. Specifically, we conduct experiments with $\mu\in\{0.4,0.6,0.8\}$ and $\sigma=1.0$. Moreover, we assume that creators’ trust degrees change dynamically over time. Let $d_{i}^{t}$ denote creator $i$’s trust degree in round $t$ and $r_{i}^{t}$ represent the number of clicks received by creator $i$ in round $t$. Then, creator $i$’s trust in round $t+1$ is updated as $d_{i}^{t+1}=d_{i}^{t}+\frac{r_{i}^{t+1}-r_{i}^{t}}{r_{i}^{t}}$, if they follow the platform’s suggestion. This means that a creator’s trust increases if following the platform’s suggestion yields more clicks, and decreases otherwise.
Figure 4 (b) shows the user clicks under trust degree distributions with different expectations on the YouTube dataset. We compare our method with Reweight, which performs best among the baselines on YouTube. The results show that although our method’s performance fluctuates slightly across different trust distributions, it consistently outperforms Reweight, demonstrating robustness to varying levels of creator trust. Furthermore, while performance of our method decreases slightly as the expected trust degree decreases, the decline is marginal. This trend is reasonable, as lower trust leads to fewer creators adopting the platform’s suggestions. These findings highlight the importance of platforms maintaining their credibility by providing beneficial suggestions to creators.
Incorporated with existing methods. We integrate our method with existing approaches that focus on recommendation algorithm design, such as P-MMF and FairRec. We exclude Reweight from this combination, as both LoRe and Reweight aim to enhance user welfare by influencing creators’ behavior. Table 3 presents the average user clicks achieved by combining our method with other approaches on the YouTube and Amazon datasets. It is evident that combining our method with these approaches further improves the performance. Moreover, LoRe + FairRec achieves better performance on the YouTube dataset, whereas LoRe + P-MMF performs better on the Amazon dataset. These results demonstrates that our approach is compatible with other methods and highlights its scalability.
6. Conclusions
In this paper, we reveal that influencing creators’ behaviors could be a better way to improve user welfare. Inspired by the idea of Bayesian persuasion, we model the platform as the sender and the creators as the receivers. To overcome the problem of bounded rationality assumption, we model the creators as bounded rational agents. Then we propose a novel method, LoRe, which aims to learn how to influence creators by revealing information to them. Lore introduces a reinforcement learning algorithm to address the limitations imposed by the strong assumption in Bayesian persuasion. Extensive experiments on two real-world datasets demonstrate the effectiveness of our approach and show that it can be seamlessly integrated into the existing RS ecosystem, as well as combined with other methods that focus on recommendation algorithm design.
Appendix
Appendix A Examples of Content Suggestion in Real Applications
Figure 6. Examples of creation suggestions in real-world applications across different short video platforms, each of which has hundreds of millions of users.
As the number of creators on short-video platforms continues to grow, providing creation suggestions to guide creators, thereby indirectly improving user experience, has become a key component of platform operations. Figure 6 presents examples of creation suggestions on short-video content platforms, each of which typically serves hundreds of millions of users.
Although current commercial platforms have begun to adopt creation suggestions, they are typically limited to simple suggestions, such as suggesting visual effects or the most trending topics. Effectively guiding creators in real-world industrial settings remains an open challenge. To bridge this gap, we aims to explore how to design more effective suggestion strategies in real-world platform environments to enhance overall user welfare.
Appendix B Omitted Definitions in Section 4
We respectively define the state value function $V^{\pi}(\bm{s}^{t})$, the state-action value function $Q^{\pi}(\bm{s}^{t},\bm{a}^{t})$as:
$V^{\pi}(\bm{s}^{t})=\operatorname{\mathbb{E}}_{(\bm{a}^{t},\bm{s}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right],$
$Q^{\pi}(\bm{s}^{t},\bm{a}^{t})=\operatorname{\mathbb{E}}_{(\bm{s}^{t+1},\bm{a}^{t+1},\dots)\sim\pi}\left[\sum_{l=0}^{\infty}\gamma^{l}\mathcal{R}(\bm{s}^{t+l},\bm{a}^{t+l},\bm{s}^{t+l+1})\right].$
Appendix C More Detailed Experimental Setups
Our experimental server is equipped with 28 Intel(R) Core(TM) i9-10940X CPUs, 128GB of RAM, and 4 NVIDIA RTX 4090 GPUs, providing a total of 96 GB of VRAM.
Appendix D Additional Experiments
D.1. Influence of Different Recommendation Models
Figure 7. Comparison of user welfare with and without our method under the RS environment with two different recommendation models, i.e, MF and BPR, respectively.
In real-world RS environments, platforms adopt different recommendation algorithms to select content from the item pool for users. To verify the effectiveness of our method under various recommendation models, we conduct experiments using different recommendation algorithms as the platform’s policy. We select two of the most widely used recommendation models: Matrix Factorization (MF) (Koren et al., 2009) and Bayesian Personalized Ranking (BPR) (Rendle et al., 2012).
As illustrated in Figure 7, across different recommendation models, our method consistently achieves better user welfare compared to the w/o signal baseline. Specifically, when using MF as the recommendation model, the user welfare with information signaling is 1.18 times that of the no-signal case (at the last round of simulation). When using BPR, this number increases to 1.26 times.
Appendix E Limitations
In our work, we use large language models (LLMs) to simulate the behaviors of users and content creators. Although previous work has verified the effectiveness of the LLM-empowered simulation agents in simulating users click behavior and creator content creation behavior, there is limited research on whether such agents can effectively simulate creators’ responses to platform-provided suggestions. Therefore, how to simulate the behavior of real-world creators after receiving platform suggestions is also a key challenge and represents a promising direction for future research.
Appendix F Social Impacts
Our method can effectively enhance the long-term user welfare in RS. This not only improves user satisfaction and engagement over time but also contributes to the development of a healthier and more sustainable ecosystem for content consumption and creation. To the best of our knowledge, our method does not pose any adverse effects on society.